OpenStudy (anonymous):

Look at the numbers below. 5, 12, 15, 16 Which pair of numbers has an odd product and an even sum?

OpenStudy (anonymous):

15 and 5

OpenStudy (saifoo.khan):

Odd product, some thing like 1,3,5 in product u have to add. Even sum = add add two terms and you should get something like, 2, 4, 6, 8

OpenStudy (anonymous):

(15,5)

OpenStudy (anonymous):

(15)(5)=75 and 15+5=20

OpenStudy (saifoo.khan):

@shayanreloaded , please dont give out direct answers. it's marked as cheating. Thanks.

OpenStudy (anonymous):

15 + 5 = 20 15 * 5 = 75

OpenStudy (anonymous):

cheating?lol i also game explanation

OpenStudy (anonymous):

He's right, why is it called cheating saifoo?

OpenStudy (saifoo.khan):

That isnt an explanation. :P

OpenStudy (saifoo.khan):

im just saying is, let the asker try first. THis way he will learn it.

OpenStudy (anonymous):

well i guess ur kinda right no spoilers....

OpenStudy (saifoo.khan):

:)